OpenStudy (anonymous):

A ball is dropped from a height of 20 feet and rebounds one fourth the distance fallen. How high does it rebound the 6th time?

OpenStudy (anonymous):

Formula r=t1/tn-1 tn=t1r^n-1

OpenStudy (anonymous):

1st rebound: 1/4 * 20 = 5 feet 2nd rebound = 1/4 * 5 = 5/4 feet and so on with the pattern (1/4)^n * 20 so 6th rebound = (1/4)^6 * 20 = 1/4096 * 20 =5/1024 feet = 0.098 feet

OpenStudy (anonymous):

nice